2. You have an order for an 8 gallon aquarium that is 20 inches long and 10.5 inches wide. How deep should the aquarium be?
this is what I know....
Common aquarium holds 10 gallons. I know volume of a right rectangular tank is found by multiplying the length of the tank times the width times the height, or V=lwh. I know 231 cubic inches are equal to 1 gallon.
Can you help me understand how to solve this? I have work due on stuff similar. Thanks!

Answer by Alan3354(69443)

You have an order for an 8 gallon aquarium that is 20 inches long and 10.5 inches wide. How deep should the aquarium be?
this is what I know....
Common aquarium holds 10 gallons. I know volume of a right rectangular tank is found by multiplying the length of the tank times the width times the height, or V=lwh. I know 231 cubic inches are equal to 1 gallon.
-----------------
What's common is not relevant.
V = lwh & 1 gallon = 231 cu inches is all you need to know.
----
V = 20*10.5*h = 8 gal*231 cubic in/gallon
210h = 1848
h = 1848/210
h =~ 8.8 inches